OpenStudy (anonymous):

*****medals and fans*****Gavin is building a fence. He needs to put a post every feet. If a new segment of the fence has 9 new posts, using the last post from the previous section of fence as a starting point, how long is the new segment of fence?

OpenStudy (anonymous):

here

OpenStudy (shinalcantara):

all you have to do is get the value of x \[\frac{ 1 }{ 9 }x = 4 \frac{ 1 }{ 2 }\] \[x = 9(4 \frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

WRONG

OpenStudy (anonymous):

its 40 1/2

OpenStudy (shinalcantara):

yep